# Improving Expressivity of Graph Neural Networks using Localization

### v3

@@ 3 Weisfeiler Leman Algorithm
Cai et al. (1992) Let $G$ and $H$ be two graphs. Then $G\simeq_{k}H$ if and only if the Duplicator wins the pebble game $BP_{k+1}(G,H).$
@@
Kiefer (2020) Let $G$ and $H$ be graphs of same size. The vertices may or may not be colored. Let $\vec{u}:=(u_{1},...,u_{k})\in(V(G))^{k}$ and $\vec{v}:=(v_{1},...,v_{k})\in(V(H))^{k}$ be any two arbitrary elements. Then, for all $i\in\mathbb{N}$, the following are equivalent :
@@ 4 Local k-WL based Algorithms for GNNs
In this section, we present the local $k-$WL based algorithms for GNNs. We also give runtime and space requirements for such GNNs.
@@ 4.1 Local k-WL
Given a graph $G$, we extract the subgraph induced on a $r$-hop neighbourhood around every vertex. We refer to it as $G_{v}^{r},$ for the subgraph rooted at vertex $v$ in $G$. Then, we colour the vertices in $G_{v}^{r}$ according to their distances from $v$. Now, we run $k-$WL on the coloured subgraph $G_{v}^{r}$. The stable colour obtained after running $k-$WL is taken as the attributes of vertex $v$. Then, we run a GNN on the graph $G$ with the attributes on each vertex $v$. This is described in Algorithm 1.
@@
4:     Color the vertices whose distance from $v$ is $i,$ by color $i$
@@
7:Each vertex has as an attribute the stable coloring of vertex $v$ obtained from $G_{v}^{r}$.
@@
Runtime and Space requirement Analysis : The time required to run $k-$WL on $n$ vertices is $O(n^{k+1}\log(n))$. Here, we run $k-$WL on a $r$-hop neighborhood instead of the entire graph. So, $n$ is replaced by $n_{1},$ where $n_{1}$ is the size of the neighborhood. If a graph has bounded degree $d$, and we run $k-$WL for a $2$-hop neighborhood, then $n_{1}$ is $O(d^{2})$. Also, we have to run Local $k-$WL for each vertex. Hence, the total time required is $O(n\cdot d^{2k+2}\log(d))$. Also, running a traditional GNN takes time $O((n+m)\log n),$ where $m$ is the number of edges. So, if we assume that $d$ is bounded, then the time required is linear in the size of the graph. Furthermore, the space required to run $k-$WL on $n$ vertices graph is $O(n^{k})$. Hence, for Local $k-$WL, it follows that the space requirement is $O(n_{1}^{k})$.
@@ 4.2 Layer k-WL
In order to make Local $k-$WL more time and space efficient, while maintaining the same expressive power, we propose a modification to Local $k-$WL. Instead of running $k-$WL on the entire $r$-hop neighbourhood, we run $k-$WL on consecutive layers of $G_{v}^{r}$ (i.e., run $k-$WL on the set of vertices with colour $i$ and colour $(i+1)$). Initially, we run $k-$WL on the set of vertices that are at a distance of $1$ and $2$ from $v.$ Then, we run $k-$WL on the set of vertices with colors $2$ and $3,$ and so on. While running $k-$WL, initially, it partitions the $k$-tuples based on the isomorphism type. However, in this setting, we incorporate the stabilized colouring obtained in the previous round. For $l<k,$ we define the color of $l$ tuples as $col(u_{1},u_{2},...,u_{l}):=col(u_{1},u_{2},...,u_{l},\underbrace{u_{1},u_{1}..,,u_{1}}_{(k-l)\text{times}})$. Consider the mixed tuple (we call a tuple to be mixed if some of the vertices have been processed in the previous iteration and the remaining have not yet been processed) $(u_{1},v_{1},\ldots,u_{k})$ where $col(u_{j})=i$ and $col(v_{j})=i+1$ (i.e $u_{i}^{\prime}$s are the set of processed vertices and $v_{i}^{\prime}$s are yet to be processed). So, even if $(u_{1},v_{1},\ldots,u_{k})$ and $(u_{1}^{{}^{\prime}},v_{1}^{{}^{\prime}},\ldots,u_{k}^{{}^{\prime}})$ may be isomorphic, if $col(u_{1},u_{2},\ldots u_{l})\neq col(u_{1}^{{}^{\prime}},u_{2}^{{}^{\prime}},\ldots u_{l}^{{}^{\prime}})$ then $col(u_{1},v_{1},\ldots,u_{k})\neq col(u_{1}^{{}^{\prime}},v_{1}^{{}^{\prime}},\ldots,u_{k}^{{}^{\prime}})$. The algorithm is described in Algorithm 2. A GNN model incorporating Local+Layer $k-$WL is equivalent to running Layer $k-$WL in line 5 in Algorithm 1.
@@
2:Run $k-$WL on the induced subgraph of levels $1$ and $2$.
@@
4:     Initial colour of $k-tuple$ incorporates the stabilized colour obtained from the previous iteration.
@@ 4.3 Recursive WL
Here, we present another variant of WL. The central idea is to decompose the graphs initially by running $1-$WL. Then, further, decompose the graphs by running $2-$WL and so on. One can note that the final vertex partition that $1-$WL outputs after color refinement are regular if we restrict to a single color class. In other words, let $G[X]$ be the induced graph on the vertices of same color. Then, $G[X]$ is regular. Also, $G[X,Y]$ where $X$ and $Y$ are sets of vertices of two different color classes. $G[X,Y]$ is also a bi-regular graph. We run $2-$WL on the regular graph. Using Bollobás (1982), we can guarantee that it would distinguish almost all regular graphs. Similarly, running $2-$WL on $G[X,Y]$ is bi-regular and thus can be distinguished by $2-$WL. We again run $1-$WL on $G,$ using the colors obtained after running $2-$WL. This further refines the colors of the vertices in $G$. One can easily check that it is more expressive than $1-$WL and less expressive than $2-$WL. We give the graph Figure 1 that can not be distinguished by Recursive $(1,2)-$WL and $1-$WL but can be distinguished by $2-$WL. This gives an intermediate hierarchy in the $k-$WL hierarchy. Also, the space and time required for running $2-$WL on the entire graph is more than that of Recursive $(1,2)-$WL. The running time and space required depend on the partition size obtained after running $1-$WL.
Note that the color of vertex $v$ is $col(v,v)$ after running $2-$WL.
@@
2:Run $1-$WL and get the partition of vertices into colour classes.
3:Let $S=\{C_{1},C_{2},\ldots C_{l}\}$ be the color classes obtained after running $1-$WL.
@@
5:     Run $2-$WL on the induced subgraph in $C_{i}$ and get color partition.
@@
8:Run $1-$WL on the colored graph $G$ whose colors are given by steps 5 and 6.
@@
10:     Run $2-$WL on the induced subgraph on the vertices in color partitions $C^{\prime}_{i}$ and $C^{\prime}_{j}$ and get new color partition.
@@
This idea can be generalized for any suitable $k.$ We can run a smaller dimensional $k_{1}-$WL and then use the partition of $k_{1}$ tuple vertices. Later, we can use this partition to get a finer partition of $k_{2}$ tuples. Assuming $k_{1}<k_{2}$, one can see that we have to run $k_{2}-$WL on smaller graphs. This reduces the time and space required for running $k_{2}-$WL on the entire graph. One can easily see that it is less expressive than $k_{2}-$WL, however, more expressive than $k_{1}-$WL.
More specifically, initially run $1-$WL and further run $(k-1)-$WL on the colored classes. One can check that it is more expressive than $(k-1)-$WL and less expressive than $k-$WL.
Figure 1: Graph identifiable by 2-WL but not by Recursive $1-$WL
@@ 5 Theoretical Guarantee of Expressive Power
In this section, we theoretically prove the expressive power of GNN models that we proposed in Section 4 in terms of graph and subgraph isomorphism. In the discussion below, we say that a GNN model $A$ is at most as expressive as a GNN model $B$ if any pair of non-isomorphic graphs $G$ and $H$ that can be distinguished by $A$ can also be distinguished by $B$. Also, we say a GNN model $A$ is at least as expressive as a GNN model $B$ if $A$ can identify all the non-isomorphic graph pairs that can be identified by GNN model $B$. The proof of the theorem and lemmas presented in the section mainly use pebble bijective game. Also, as mentioned earlier, there is equivalence between the expressivity of $k-$WL and $(k+1)$-pebble bijective game.
@@ 5.1 Local k-WL
It has been shown in recent works that running Local $1-$WL has more expressive power as compared to running $1-$WL. The upper bound on the expressive power of Local $1-$WL has been shown in Frasca et al. (2022). However, the expressive power of Local $k-$WL, for arbitrary $k,$ has not been studied. In the Theorem 3, we give a description of the expressive power of Local $k-$WL and show that it has more expressive power than $k-$WL. We also show that it is less expressive than $(k+1)-$WL. The proof techniques that we used are different from Frasca et al. (2022).
@@
Let $G_{1}$ and $G_{2}$ be two graphs distinguishable by $k-$WL. So, the Spoiler has a winning strategy in the game ($G_{1}$,$G_{2}$). Suppose $G_{1}$ and $G_{2}$ do not get distinguished after running $k-$WL locally. That means the Duplicator has a winning strategy for all vertices individualized. Let $v$ in $G_{1}$ and $u$ in $G_{2}$ be the vertices that are individualized.
@@
Initially, $D_{2}$ gives a bijection $f$ from the vertex set of $G_{1}^{v}$ to $G_{2}^{u}$. We propose the same bijection $f$ by $D_{1},$ extending it by mapping $v$ to $u$. Now, the spoiler $S_{1}$ places a pebble at some vertex $(v_{i},f(v_{i})).$ The spoiler $S_{2}$ also places a pebble at vertex $(v_{i},f(v_{i}))$. We can show using induction on the number of rounds that if $S_{1}$ wins the game, then $S_{2}$ also wins the game. Our induction hypothesis is that $S_{1}$ has not won till the $j^{th}$ round and the positions of both the games are same. Let the current position of both the games after the $j^{th}$ round be $((v_{1},v_{2},\ldots,v_{l}),(f(v_{1}),f(v_{2}),\ldots,f(v_{l}))$. Now, $S_{1}$ either decides to play a pair of pebbles or remove.
@@
Case(2): If $S_{1}$ decides to play a pebble. In this case, $S_{2}$ also decides to play a pebble. Duplicator $D_{2}$ proposes a bijective function $f.$ The same bijective function is proposed by $D_{1}$. Now, $S_{1}$ places a pebble at $(v,f(v))$. $S_{2}$ also chooses the same vertices. So, the position of both the game is the same. Therefore, if $S_{1}$ wins the game, then $S_{2}$ also wins the game. Thus, running $k-$WL locally is at least as expressive as running $k-$WL on the entire graph.
We can show that it is more expressive by looking at the simple example that running $1-$WL on a local substructure can count the number of triangles, whereas running $1-$WL on an entire graph does not recognize graphs having different triangle counts. Also, one can observe that running $k-$WL locally is running $k-$WL on the colored graph where vertices at distinct distances get distinct colors. Its power is the same as individualizing one vertex and running $k-$WL. Thus, running $k-$WL locally is more expressive than running $k-$WL on the entire graph.
Let $G_{1}$ and $G_{2}$ be two graphs that can be distinguished by running $k-$WL locally. Recall that, the key vertices refer to $u$ and $v$ in $G_{1}$ and $G_{2}$ such that they are the root vertices corresponding to $G_{1}$ and $G_{2}$, respectively. This means the Spoiler has a winning strategy in the $(k+1)$ bijective pebble game, where the key vertices are matched to each other. Now, we use the strategy of the Spoiler in the local substructure to get a winning strategy for the Spoiler in the entire graph. At first, when the Duplicator gives a bijective function, the Spoiler places a pebble on the paired vertices. For the remaining moves, we copy the strategy of the Spoiler in the local structure, and the Duplicator’s strategy of the entire graph is copied to the Duplicator’s strategy of the local structures. Thus, if the Spoiler has a winning strategy in the local substructure, then the Spoiler wins the $(k+2)-$ bijective pebble game on entire graphs. ∎
@@ 5.2 Layer k-WL
Let $G$ and $H$ be the subgraphs induced on the $r$-hop neighborhood. Let $(S,D)$ be the Spoiler-Duplicator pair for the game $(G,H)$. Similarly, let $(S_{i},D_{i})$ be the Spoiler-Duplicator pair for the game $(G_{i},H_{i}),$ where $G_{i}$ and $H_{i}$ are the subgraphs induced on the vertices at the $i$th and $(i+1)$th layers of $G$ and $H,$ respectively. We claim that if any of the $S_{i}^{\prime}$s has a winning strategy in the game $(G_{i},H_{i})$, then $S$ has a winning strategy in the game $(G,H)$. Here, the strategy of $D$ is copied by $D_{i}$, and the strategy of $S_{i}$ is copied by $S$. We prove this using induction on the number of rounds of the game. Our induction hypothesis is that the positions of both the games are same, and if $S_{i}$ wins after $t$ rounds, then $S$ also wins after $t$ rounds.
Base case: $D$ proposes a bijective function $f:V(G)\xrightarrow{}V(H)$. Note that the bijection must be color-preserving; otherwise, $S$ wins in the first round. Thus, we can assume that $f$ is color-preserving. So, $D_{i}$ proposes the restricted function $f_{i}$ as a bijective function from $V(G_{i})$ to $V(H_{i})$. Now, $S_{i}$ plays a pair of pebbles in $(G_{i},H_{i})$, and $S$ also plays the same pair of pebbles in the game $(G,H)$. It is easy to see that both games’ positions are the same. Also, if $S_{i}$ wins, then the number of vertices of a particular color is different. Hence, $S$ also has a winning strategy.
@@
Consider the $(t+1)^{th}$ round in both games. $S_{i}$ either chooses to play or remove a pebble. If $S_{i}$ chooses to remove a pebble, so will $S$. Again, the position of both the games is same. Now, if $S_{i}$ decides to play a pair of pebbles, then $S$ also decides to play a pair of pebbles. So, $D$ proposes a bijective function and $D_{i}$ proposes a restricted bijective function. Now, if $S_{i}$ plays a pair of pebbles at $(v,f_{i}(v))$, then $S$ also decides to play a pair of pebbles at $(v,f(v))$. Thus, the position of the game is same in both of the games. This ensures that if $S_{i}$ has won, then $S$ also wins.
@@
Let $G$ and $H$ be the subgraphs induced on a $r$-hop neighborhood. Let $(S,D)$ be the Spoiler-Duplicator pair for the game $(G,H)$. Similarly, let $(S_{i},D_{i})$ be the Spoiler-Duplicator pair for the game $(G_{i},H_{i}),$ where $G_{i}$ and $H_{i}$ are the subgraphs induced on the vertices at the $i$th and $(i+1)$th layers of $G$ and $H,$ respectively. We claim that if $S$ has a winning strategy in the game $(G,H),$ then there exists $S_{i}$ such that it has a winning strategy in the game $(G_{i},H_{i}).$ Here, the strategy of $D$ is copied by $D_{i}$ and the strategy of $S_{i}$ is copied by $S.$ We prove the lemma using induction on the number of rounds of the game. Our induction hypothesis is that the position of the game $(G,H)$ is same for $(G_{i},H_{i}),$ for all $i,$ if we restrict it to the subgraph induced by the vertices of color $i$ and $(i+1).$ Also, if $S$ wins after round $t,$ then there exists $S_{i}$ that wins after $t$ rounds.
@@
Base case: $D$ proposes a bijective function $f:V(G)\longrightarrow V(H)$. Note that the bijection must be color-preserving; otherwise $S$ wins in the first round. Thus, we can assume that $f$ is color-preserving. So, $D_{i}$ proposes the restricted function $f_{i}$ as a bijective function from $V(G_{i})$ to $V(H_{i}),\forall i\in[r]$. Now, $S$ will play a pair of pebbles in the game $(G,H).$ Suppose $S$ plays the pebbles at $(v,f(v))$ and $color(v)=i,$ then $S_{i}$ and $S_{i-1}$ play pebbles at $(v,f_{i}(v))$ in their first round. It is easy to see that the position of the games $(G,H)$ and $(G_{i},H_{i}),$ for all $i\in[r],$ is same if we restrict it to the subgraph induced by vertices of colors $i$ and $(i+1)$. Also, if $S$ wins, then the number of vertices of particular color are not same. So, there exists some $i,$ such that $S_{i}$ also has a winning strategy.
@@
By induction hypothesis, assume that after the $t^{th}$ round, $S$ did not win and position of the game is same as defined.

### v5

@@ 3 Weisfeiler Leman Algorithm
(cai1992optimal) Let $G$ and $H$ be two graphs. Then $G\simeq_{k}H$ if and only if the Duplicator wins the pebble game $BP_{k+1}(G,H).$
@@
(kiefer2020power) Let $G$ and $H$ be graphs of the same size. The vertices may or may not be coloured. Let $\vec{u}:=(u_{1},\dots,u_{k})\in(V(G))^{k}$ and $\vec{v}:=(v_{1},\dots,v_{k})\in(V(H))^{k}$ be any two arbitrary elements. Then, for all $i\in\mathbb{N}$, the following are equivalent :
@@ 4 Local k-WL based Algorithms for GNNs
This section presents the local $k$-WL based algorithms for GNNs. We also give runtime and space requirements for such GNNs.
@@ 4.1 Local k-WL
Given a graph $G$, we extract the subgraph induced on a $r$-hop neighbourhood around every vertex. We refer to it as $G_{v}^{r},$ for the subgraph rooted at vertex $v$ in $G$. Then, we colour the vertices in $G_{v}^{r}$ according to their distances from $v$. Now, we run $k$-WL on the coloured subgraph $G_{v}^{r}$. The stable colour obtained after running $k$-WL is taken as the attributes of vertex $v$. Then, we run a GNN on the graph $G$ with the attributes on each vertex $v$. This is described in 1.
@@
4:  Color the vertices whose distance from $v$ is $i,$ by color $i$.
@@
7:Each vertex has as an attribute the stable colouring of vertex $v$ obtained from $G_{v}^{r}$.
@@
Runtime and Space requirement Analysis : The time required to run $k$-WL on $n$ vertices is $O(n^{k+1}\log(n))$. Here, we run $k$-WL on a $r$-hop neighbourhood instead of the entire graph. So, $n$ is replaced by $n_{1},$ where $n_{1}$ is the size of the neighbourhood. If a graph has bounded degree $d$, and we run $k$-WL for a 22-hop neighbourhood, then $n_{1}$ is $O(d^{2})$. Also, we have to run Local $k$-WL for each vertex. Hence, the total time required is $O(n\cdot d^{2k+2}\log(d))$. Also, running a traditional GNN takes time $O((n+m)\log n),$ where $m$ is the number of edges. So, if we assume that $d$ is bounded, then the time required is linear in the size of the graph. Furthermore, the space required to run $k$-WL on $n$ vertices graph is $O(n^{k})$. Hence, for Local $k$-WL, it follows that the space requirement is $O(n_{1}^{k})$.
@@ 4.2 Layer k-WL
In order to make Local $k$-WL more time and space-efficient while maintaining the same expressive power, we propose a modification to Local $k$-WL. Instead of running $k$-WL on the entire $r$-hop neighbourhood, we run $k$-WL on consecutive layers of $G_{v}^{r}$ (i.e., run $k$-WL on the set of vertices with colour $i$ and colour $(i+1)$). Initially, we run $k$-WL on the set of vertices that are at a distance of $1$ and $2$ from $v.$ Then, we run $k$-WL on the set of vertices with colors $2$ and $3,$ and so on. While running $k$-WL, initially, it partitions the $k$-tuples based on the isomorphism type. However, we incorporate the stabilized colouring obtained in the previous round in this setting. For $l<k,$ we define the color of $l$ tuples as $col(u_{1},u_{2},...,u_{l}):=col(u_{1},u_{2},...,u_{l},\underbrace{u_{1},u_{1}..,,u_{1}}_{(k-l)\text{times}})$. Consider the mixed tuple (we call a tuple to be mixed if some of the vertices have been processed in the previous iteration and the remaining have not yet been processed) $(u_{1},v_{1},\ldots,u_{k})$ where $col(u_{j})=i$ and $col(v_{j})=i+1$ (i.e $u_{i}^{\prime}$s are the set of processed vertices and $v_{i}^{\prime}$s are yet to be processed). So, even if $(u_{1},v_{1},\ldots,u_{k})$ and $(u_{1}^{{}^{\prime}},v_{1}^{{}^{\prime}},\ldots,u_{k}^{{}^{\prime}})$ may be isomorphic, if $col(u_{1},u_{2},\ldots u_{l})\neq col(u_{1}^{{}^{\prime}},u_{2}^{{}^{\prime}},\ldots u_{l}^{{}^{\prime}})$ then $col(u_{1},v_{1},\ldots,u_{k})\neq col(u_{1}^{{}^{\prime}},v_{1}^{{}^{\prime}},\ldots,u_{k}^{{}^{\prime}})$. The algorithm is described in 2. A GNN model incorporating Local+Layer $k$-WL is equivalent to running Layer $k$-WL in line 5 in 1.
@@
2:Run $k$-WL on the induced subgraph of levels 11 and $2$.
@@
4:  Initial colour of ktuplek-tuple incorporates the stabilized colour obtained from the previous iteration.
@@ 4.3 Recursive WL
Here, we present another variant of WL. The central idea is to decompose the graphs initially by running 11-WL. Then, further, decompose the graphs by running $2$-WL and so on. One can note that the final vertex partition that 11-WL outputs after colour refinement is regular if we restrict it to a single colour class. In other words, let $G[X]$ be the induced graph on the vertices of the same colour. Then, $G[X]$ is regular. Also, $G[X,Y]$ where $X$ and $Y$ are sets of vertices of two different colour classes. $G[X,Y]$ is also a bi-regular graph. We run $2$-WL on the regular graph. Using (bollobas1982distinguishing), we can guarantee that it would distinguish almost all regular graphs. Similarly, running $2$-WL on $G[X,Y]$ is bi-regular and thus can be distinguished by 22-WL. We again run $1$-WL on $G,$ using the colours obtained after running 22-WL. This further refines the colours of the vertices in $G$. One can easily check that it is more expressive than $1$-WL and less expressive than $2$-WL. We give the graph 1 that can not be distinguished by Recursive $(1,2)$-WL and 11-WL but can be distinguished by 22-WL. This gives an intermediate hierarchy in the $k$-WL hierarchy. Also, the space and time required for running $2$-WL on the entire graph is more than that of Recursive $(1,2)$-WL. The running time and space required depend on the partition size obtained after running 11-WL.
Note that the colour of vertex $v$ is $col(v,v)$ after running 22-WL.
@@
2:Run 11-WL and get the partition of vertices into colour classes.
3:Let $S=\{C_{1},C_{2},\ldots C_{l}\}$ be the color classes obtained after running 11-WL.
@@
5:  Run 22-WL on the induced subgraph in $C_{i}$ and get color partition.
@@
8:Run 11-WL on the colored graph $G$ whose colors are given by steps 5 and 6.
@@
10:  Run 22-WL on the induced subgraph on the vertices in color partitions $C^{\prime}_{i}$ and $C^{\prime}_{j}$ and get new color partition.
@@
This idea can be generalized for any suitable $k.$ We can run a smaller dimensional $k_{1}$-WL and then use the partition of $k_{1}$ tuple vertices. Later, we can use this partition to get a finer partition of $k_{2}$ tuples. Assuming $k_{1}<k_{2}$, one can see that we have to run $k_{2}$-WL on smaller graphs. This reduces the time and space required for running $k_{2}$-WL on the entire graph. One can easily see that it is less expressive than $k_{2}$-WL; however, it is more expressive than $k_{1}$-WL.
More specifically, initially run $1$-WL and further run (k1)(k-1)-WL on the coloured classes. One can check that it is more expressive than (k1)(k-1)-WL and less expressive than $k$-WL.
Figure 1: Graphs identifiable by 2-WL but not by Recursive 11-WL
@@ 5 Theoretical Guarantee of Expressive Power
In this section, we theoretically prove the expressive power of GNN models that we proposed in 4 in terms of graph and subgraph isomorphism. In the discussion below, we say that a GNN model $A$ is at most as expressive as a GNN model $B$ if any pair of non-isomorphic graphs $G$ and $H$ that can be distinguished by $A$ can also be distinguished by $B$. Also, we say a GNN model $A$ is at least as expressive as a GNN model $B$ if $A$ can identify all the non-isomorphic graph pairs that can be identified by GNN model $B$. The proof of the theorem and lemmas presented in the section mainly use pebble bijective game. Also, as mentioned earlier, there is equivalence between the expressivity of $k$-WL and $(k+1)$-pebble bijective game.
@@ 5.1 Local k-WL
It has been shown in recent works that running Local $1$-WL has more expressive power as compared to running 11-WL. The upper bound on the expressive power of Local $1$-WL has been shown in (frasca). However, the expressive power of Local $k$-WL, for arbitrary $k,$ has not been studied. In the Theorem 3, we give a description of the expressive power of Local $k$-WL and show that it has more expressive power than $k$-WL. We also show that it is less expressive than $(k+1)$-WL. The proof techniques that we used are different from (frasca).
@@
Let $G_{1}$ and $G_{2}$ be two graphs distinguishable by $k$-WL. So, the Spoiler has a winning strategy in the game ($G_{1}$$G_{2}$). Suppose $G_{1}$ and $G_{2}$ do not get distinguished after running $k$-WL locally. That means the Duplicator has a winning strategy for all vertices individualized. Let $v$ in $G_{1}$ and $u$ in $G_{2}$ be the vertices that are individualized.
@@
Initially, $D_{2}$ gives a bijection $f$ from the vertex set of $G_{1}^{v}$ to $G_{2}^{u}$. We propose the same bijection $f$ by $D_{1},$ extending it by mapping $v$ to $u$. Now, the spoiler $S_{1}$ places a pebble at some vertex $(v_{i},f(v_{i})).$ The spoiler $S_{2}$ also places a pebble at vertex $(v_{i},f(v_{i}))$. We can show using induction on the number of rounds that if $S_{1}$ wins the game, then $S_{2}$ also wins the game. Our induction hypothesis is that $S_{1}$ has not won till the $j^{th}$ round, and the positions of both games are the same. Let the current position of both the games after the $j^{th}$ round be $((v_{1},v_{2},\ldots,v_{l}),(f(v_{1}),f(v_{2}),\ldots,f(v_{l}))$. Now, $S_{1}$ either decides to play a pair of pebbles or remove.
@@
Case(2): If $S_{1}$ decides to play a pebble. In this case, $S_{2}$ also decides to play a pebble. Duplicator $D_{2}$ proposes a bijective function $f.$ The same bijective function is proposed by $D_{1}$. Now, $S_{1}$ places a pebble at $(v,f(v))$. $S_{2}$ also chooses the same vertices. So, the position of both the games is the same. Therefore, if $S_{1}$ wins the game, then $S_{2}$ also wins the game. Thus, running $k$-WL locally is at least as expressive as running $k$-WL on the entire graph.
We can show that it is more expressive by looking at the simple example that running $1$-WL on a local substructure can count the number of triangles, whereas running $1$-WL on an entire graph does not recognize graphs having different triangle counts. Also, one can observe that running $k$-WL locally is running $k$-WL on the coloured graph where vertices at distinct distances get distinct colours. Its power is the same as individualizing one vertex and running $k$-WL. Thus, running $k$-WL locally is more expressive than running $k$-WL on the entire graph.
Let $G_{1}$ and $G_{2}$ be two graphs that can be distinguished by running $k$-WL locally. Recall that the key vertices refer to $u$ and $v$ in $G_{1}$ and $G_{2}$ such that they are the root vertices corresponding to $G_{1}$ and $G_{2}$, respectively. This means the Spoiler has a winning strategy in the $(k+1)$ bijective pebble game, where the key vertices are matched to each other. Now, we use the strategy of the Spoiler in the local substructure to get a winning strategy for the Spoiler in the entire graph. At first, when the Duplicator gives a bijective function, the Spoiler places a pebble on the paired vertices. For the remaining moves, we copy the strategy of the Spoiler in the local structure, and the Duplicator’s strategy of the entire graph is copied to the Duplicator’s strategy of the local structures. Thus, if the Spoiler has a winning strategy in the local substructure, then the Spoiler wins the $(k+2)$- bijective pebble game on entire graphs. ∎
@@ 5.2 Layer k-WL
Let $G$ and $H$ be the subgraphs induced on the $r$-hop neighbourhood. Let $(S,D)$ be the Spoiler-Duplicator pair for the game $(G,H)$. Similarly, let $(S_{i},D_{i})$ be the Spoiler-Duplicator pair for the game $(G_{i},H_{i}),$ where $G_{i}$ and $H_{i}$ are the subgraphs induced on the vertices at the $i$th and $(i+1)$th layers of $G$ and $H,$ respectively. We claim that if any of the $S_{i}^{\prime}$s has a winning strategy in the game $(G_{i},H_{i})$, then $S$ has a winning strategy in the game $(G,H)$. Here, the strategy of $D$ is copied by $D_{i}$, and the strategy of $S_{i}$ is copied by $S$. We prove this using induction on the number of rounds of the game. Our induction hypothesis is that the positions of both the games are the same, and if $S_{i}$ wins after $t$ rounds, then $S$ also wins after $t$ rounds.
Base case: $D$ proposes a bijective function $f:V(G)\xrightarrow{}V(H)$. Note that the bijection must be colour-preserving; otherwise, $S$ wins in the first round. Thus, we can assume that $f$ is color-preserving. So, $D_{i}$ proposes the restricted function $f_{i}$ as a bijective function from $V(G_{i})$ to $V(H_{i})$. Now, $S_{i}$ plays a pair of pebbles in $(G_{i},H_{i})$, and $S$ also plays the same pair of pebbles in the game $(G,H)$. It is easy to see that the positions of the two games are the same. Also, if $S_{i}$ wins, then the number of vertices of a particular colour is different. Hence, $S$ also has a winning strategy.
@@
Consider the $(t+1)^{th}$ round in both games. $S_{i}$ either chooses to play or remove a pebble. If $S_{i}$ chooses to remove a pebble, so will $S$. Again, the position of both the games is the same. If $S_{i}$ decides to play a pair of pebbles, then $S$ also decides to play a pair of pebbles. So, $D$ proposes a bijective function, and $D_{i}$ proposes a restricted bijective function. Now, if $S_{i}$ plays a pair of pebbles at $(v,f_{i}(v))$, then $S$ also decides to play a pair of pebbles at $(v,f(v))$. Thus, the position of the game is the same in both of the games. This ensures that if $S_{i}$ has won, then $S$ also wins.
@@
Let $G$ and $H$ be the subgraphs induced on a $r$-hop neighbourhood. Let $(S,D)$ be the Spoiler-Duplicator pair for the game $(G,H)$. Similarly, let $(S_{i},D_{i})$ be the Spoiler-Duplicator pair for the game $(G_{i},H_{i}),$ where $G_{i}$ and $H_{i}$ are the subgraphs induced on the vertices at the $i$th and $(i+1)$th layers of $G$ and $H,$ respectively. We claim that if $S$ has a winning strategy in the game $(G,H),$ then there exists $S_{i}$ such that it has a winning strategy in the game $(G_{i},H_{i}).$ Here, the strategy of $D$ is copied by $D_{i}$ and the strategy of $S_{i}$ is copied by $S.$ We prove the lemma using induction on the number of rounds of the game. Our induction hypothesis is that the position of the game $(G,H)$ is the same for $(G_{i},H_{i}),$ for all $i,$ if we restrict it to the subgraph induced by the vertices of color $i$ and $(i+1).$ Also, if $S$ wins after round $t,$ then there exists $S_{i}$ that wins after $t$ rounds.
@@
Base case: $D$ proposes a bijective function $f:V(G)\longrightarrow V(H)$. Note that the bijection must be colour-preserving; otherwise, $S$ wins in the first round. Thus, we can assume that $f$ is color-preserving. So, $D_{i}$ proposes the restricted function $f_{i}$ as a bijective function from $V(G_{i})$ to $V(H_{i}),\forall i\in[r]$. Now, $S$ will play a pair of pebbles in the game $(G,H).$ Suppose $S$ plays the pebbles at $(v,f(v))$ and $color(v)=i,$ then $S_{i}$ and $S_{i-1}$ play pebbles at $(v,f_{i}(v))$ in their first round. It is easy to see that the position of the games $(G,H)$ and $(G_{i},H_{i}),$ for all $i\in[r],$ is the same if we restrict it to the subgraph induced by vertices of colours $i$ and $(i+1)$. Also, if $S$ wins, then the number of vertices of a particular colour is not the same. So, there exists some $i,$ such that $S_{i}$ also has a winning strategy.
@@
By induction hypothesis, assume that after the $t^{th}$ round, $S$ did not win, and the position of the game is the same as defined.